OpenStudy (anonymous):

Help please! . If (a/b) = (c/d), which of these statements is not true? a) ac = bd, given that b and d are not 0. < - b) b/a = d/c, given that a, b, c, and d are not 0. c) ad = bc, given that b and d are not 0. d) (a + b)/b = (c + d)/d, given that b and d are not 0.

OpenStudy (anonymous):

if it is not clear, try it with numbers and see which one is false for example, put \[\frac{2}{3}=\frac{4}{6}\]i.e. make \(a=2,b=3,c=4,d=6\) and check which one of the 4 statements is false

OpenStudy (anonymous):

Okay let me do that

OpenStudy (anonymous):

i think you will get it right away

OpenStudy (anonymous):

So the first one is right. Im going to try the second one.

OpenStudy (anonymous):

hmm not sure what you mean by "the first one is right"

OpenStudy (anonymous):

I mean it works with what its saying. Like number 2 is also correct, it does nto equal 0

OpenStudy (anonymous):

unless you mean "right, it is NOT true" since \(2\times 4\neq 3\times 6\)

OpenStudy (anonymous):

Right. So would C be the correct answer?

OpenStudy (anonymous):

i seem to have lost you you have to pick which one is NOT true

OpenStudy (anonymous):

OHhhhhhh. Gah. Okay. Im silly. Im sorry.

OpenStudy (anonymous):

clear right? the first one is NOT true, so that one is your answer

OpenStudy (anonymous):

Okay. Thank you! That makes more sense. I was misreading the question